a(cosA+cosB)=2(a+b)sin^2A/2

Grade:11

4 Answers

Shaurya Gupta
11 years ago

question is wrong. It is 2(a+b)sin2(C/2). I solved it.

Shaurya Gupta
11 years ago

dude after solving this it comes out 2(a+b)[c2 -(a 2+b2 )]/4ab = 2(a+b)sin2c/2

if u wanna try just put the values of cosA and cosB and then take the LCM and solve.

Shaurya Gupta
11 years ago

Well the answer is coming 2(a+b)(c^2-(a-b)^2)/4ab = 2(a+b)sin^2(C/2)

Take the LCM of values of cosa and  cosb and then solve. it will come this only

orange panda
11 years ago

the question might be wrong. pleez check again